# Scalable analysis of stop-and-go waves: Representation, measurements and insights

### v2

@@ 2.2 Stop-and-go waves identification techniques
On the identification method, the central challenge lies in identifying the areas of interest across both spatial and temporal dimensions. Table 2 summarizes the method used to identify the stop-and-go waves. Using speed contour lines [4; 59] is the most straightforward approach for analyzing stop-and-go waves; however, it is inadequate for distinguishing the individual structures of each wave. Time-series data analysis [77; 94] within the time domain for each spatial or vehicle observation is another method to measure stop-and-go waves. Another mainstream method is to transform the data to the frequency domain, which allows for the separation of low-frequency and high-frequency signals [19; 20; 33], where the Wavelet Transform is viewed as the best of practice. However, all of the aforementioned methods primarily focus on wave generation and propagation, often neglecting the phenomena of wave merging and bifurcation. [16] provided a detailed behavioral analysis using asymmetric traffic theory, illustrating the wave propagation path across both space and time. This analysis revealed that the path of acceleration curves may split, potentially generating a secondary wave. This highlights the need for methods that not only detect and track waves but also stitch them together over time to support the full life-cycle wave analysis.
@@ 2.3 Emerging large-scale trajectory data
As stated by [7], more trajectory data are needed. In recent years, more trajectory data collection efforts comes from drones [95; 8; 39; 96; 97; 40], helicopter [40], and roadside infrastructures [9; 98; 61; 11]. As more data is generated, the need for tools to analyze this data and enhance research reproducibility within the community becomes more evident [99]. Additionally, with the continuous collection of data, the size of the datasets has become significantly larger than previous ones, adding more challenges for researchers to address [100]. Moreover, these datasets may uncover previously unobserved scenarios, such as the effects of automated control vehicles [40], traffic smoothing strategies [81; 101], electric vehicles [53], and traffic incidents [102]. While generating more data is a crucial first step, the development of more analytical tools is equally paramount. The majority of existing tools do not open-source their analysis software (shown in Table 2), which potentially limits subsequent analyses on different datasets. This paper aims to fill this gap by exploring the analysis of massive trajectory data to make stop-and-go wave analysis at scale possible.
@@ 4 Methodology
$\bm{\mathrm{\tau}}_{i}$ vector representing the trajectory of vehicle $i$ - -
$\bm{\mathrm{\tau}}_{i}^{c}$ vector representing the stationary component trajectory of vehicle $i$ - -
$\bm{\mathrm{\tau}}_{i}^{osc}$ vector representing the oscillation component trajectory of vehicle $i$ - -
@@ 4.1 Preliminaries
Long-range raw vehicle trajectory data as from [11] is imperfect due to the challenges of tracking and re-identification in dense traffic conditions [103; 104]. In essence, it is difficult to track vehicles perfectly across hundreds of cameras and through occlusions [105], so trajectory datasets over large spatial ranges often contain fragmented trajectories. These issues do not significantly impact local traffic measurements, such as traffic speeds, which can still be obtained with extremely high fidelity. We prepare raw data for the methods described next by utilizing the method from [106], which we describe in more detail in A. The resulting virtual trajectories have no passing or lane change maneuvers. This allows efficient search between adjacent trajectories, but the methods described next could also be applied to raw trajectory data with lane-change maneuvers with slight modifications.
@@ 4.2 Identifying the stop-and-go waves: the trajectory perspective
Let $\mathcal{T}$ be the set of all trajectories. Note that the preprocessing in A ensures that trajectories are ordered by increasing time, and no trajectories enter or exit the considered lane. Let $\bm{\mathrm{\tau}}_{i}\in\mathcal{T}$ represent the trajectory of vehicle $i$ as a vector of time-space points. Throughout this section, we will use $\tau_{i}(t)$ to represent a continuous, functional representation of the trajectory of vehicle $i$ from which the points in $\bm{\mathrm{\tau}}_{i}$ are sampled. To detect the critical events from the trajectories, each trajectory is decomposed into two components as follows:
@@
In our implementation, the local maxima and minima of $\tau_{i}^{\text{osc}}(t)$ are identified using peak-finding algorithms scipy.signal.find_peaks [107]. We find all local maxima and minima in $\bm{\mathrm{\tau}}_{i}^{osc}$, yielding a set $d_{i}$ of wave front and a set $a_{i}$ of wave tail points for trajectory $i$. We keep only those wave fronts that are followed by a wave tail and those wave tails that are preceded by a wave front, excluding any wave fronts or tails that occur at the boundaries of the trajectories, as they do not contribute to a complete stop-and-go cycle. This ensures that there are the same number $K_{i}$ of wave front points and wave tail points for trajectory $i$. This process is repeated for all trajectories in $\mathcal{T}$, yielding an overall set of wave front and tail points $\mathcal{D}=\{d_{0},d_{1},...d_{i}\}$ and $\mathcal{A}=\{a_{0},a_{1},...a_{i}\}$.
@@ 5 Data
I-24 MOTION is a traffic instrument for the freeway traffic observation [111; 11; 104], which is designed for the continuous freeway traffic data collection and analysis. It is expected to generate 200 million vehicle miles traveled (VMT) in trajectory data annually [104], which poses significant challenges for data analysis. In this paper, the I-24 MOTION INCEPTION data released by [11] is used for the stop-and-go wave analysis with the proposed method in Section 4. Table 4 summarizes the data used in this paper, including the corresponding week day of the date, vehicle miles traveled (VMT) and the mean speed (MS). The method input consists of virtual trajectories generated from the mean speed field, designed to address data quality issues [103] in individual trajectories and enable scalable measurements since traffic waves are the phenomena in between microscopic and macroscopic. Further details are provided in A.
@@ 7.5 Spatial distribution of wave bifurcation and merge
Although identifying the root cause of bifurcation and merging is challenging, our analysis of road geometry suggests that these phenomena can occur anywhere, although with a concentration within the ramp impact area. We believe the underlying cause is more closely tied to driving behavior, with road geometry acting as a contributing factor by influencing the frequency of cut-ins and lane changes. Factors such as asymmetric driving behavior [114; 16], anticipation effects [115], and driver distraction [73] may play a significant role. A more detailed microscopic analysis of driving behavior, enabled by improvements in data quality [103], could provide deeper insights in the future.
@@ 8 Conclusion
The code developed in this article is publicly available at https://github.com/I24-MOTION/wave-analysis and will be listed on https://i24motion.org. The tools introduced in this article could open new possibilities for analyzing stop-and-go traffic waves at scale and it could lower the barrier for researchers to deal with large-scale trajectory data. For example, the I-24 MOTION instrument [11] will enable both CAV field experiments [81; 120] and ITS deployments [121]. It is worthy of attention that the method and tool presented in this article allow for empirically analyzing how control strategies and vehicle automation influence traffic wave properties, a growing topic within traffic flow theory and characteristics [122; 123; 124; 93; 119; 120]. Further research will focus on investigating the factors influencing wave evolution [65], accompanied by refinements in data quality.
@@ A.1 Virtual sensors and virtual vehicles
Long-range raw vehicle trajectory data is imperfect due to the challenges of tracking and re-identification in dense traffic conditions [103; 104]. The tasks fail when errors occurring at the current frame cannot be corrected using information from subsequent frames. These errors may arise from various system sources, including homography estimation [103] and object occlusion [105]. However, these challenges do not significantly impact local traffic measurements, such as traffic speeds, which can still be obtained with extremely high fidelity. In this context, generating vehicle trajectories [125] from high-fidelity speed field can serve as a surrogate measure for the raw trajectory data, enabling the analysis of stop-and-go waves.We use the virtual trajectory generation method from [106] as a first step in wave identification in this work. We detail this approach below. Table 3 provides a comprehensive list of the variables and parameters used throughout the appendix.
@@ A.1.1 Virtual sensors: Edie’s definition for macroscopic measurements
Edie [126] provides an approach to calculate spatio-temporal mean of density, flow speed, and traffic flow from vehicle trajectories, which is named as Edie’s definition (shown in Figure 15). According to the definition, the density and the flow can be computed from the total travel time (TTT) and total travel distance (TTD) of all the vehicles within an area. Using Edie’s definition allows for the parallel calculation of measurements on a trajectory-by-trajectory basis [106]. Consider a shear box [127; 125] centered at a point ($t$, $x$), and let $\Delta t$ and $\Delta x$ denote the height and width of the box, as shown in Figure 15. The macroscopic estimates can be computed as:
@@ A.1.2 Adaptive Smoothing
With the original speed field generated, virtual trajectories can be created. However, the original macroscopic measurements often contain missing values and outliers. As suggested by [125], an adaptive smoothing method should be applied to the raw data to improve accuracy. The Adaptive Smoothing Method (ASM), developed by [128; 129], is a widely used smoothing and interpolation algorithm for constructing a continuous spatio-temporal mean speed field. While it is especially useful for data collected from fixed infrastructure sensors, such as inductive loops, it is also applicable to the macroscopic speed data generated from trajectories. This method is particularly helpful in interpolating speeds in occluded areas, such as under bridges. We directly apply this method, with parameter settings as listed in Table 9.
@@
$\sigma$(mile) smoothing width in time coordinate 0.12
$\tau$(second) smoothing width in space coordinate 20

### v3

@@ 2.2 Stop-and-go waves identification techniques
On the identification method, the central challenge lies in identifying the areas of interest across both spatial and temporal dimensions. Table 2 summarizes the method used to identify the stop-and-go waves. Using speed contour lines [4; 59] is the most straightforward approach for analyzing stop-and-go waves; however, it is inadequate for distinguishing the individual structures of each wave. Time-series data analysis [77; 94] within the time domain for each spatial or vehicle observation is another method to measure stop-and-go waves. Another mainstream method is to transform the data to the frequency domain, which allows for the separation of low-frequency and high-frequency signals [19; 20; 33], where the Wavelet Transform is viewed as the best of practice. More recent studies [95] leverage graph-based representations to provide insights into congestion patterns. Moreover, the aforementioned methods primarily focus on wave generation and propagation, often neglecting the phenomena of wave merging and bifurcation. [16] provided a detailed behavioral analysis using asymmetric traffic theory, illustrating the wave propagation path across both space and time. This analysis revealed that the path of acceleration curves may split, potentially generating a secondary wave. This highlights the need for methods that not only detect and track waves but also stitch them together over time to support the full life-cycle wave analysis.
@@ 2.3 Emerging large-scale trajectory data
As stated by [7], more trajectory data are needed. In recent years, more trajectory data collection efforts comes from drones [96; 8; 39; 97; 98; 40], helicopter [40], and roadside infrastructures [9; 99; 61; 11]. As more data is generated, the need for tools to analyze this data and enhance research reproducibility within the community becomes more evident [100]. Additionally, with the continuous collection of data, the size of the datasets has become significantly larger than previous ones, adding more challenges for researchers to address [101]. Moreover, these datasets may uncover previously unobserved scenarios, such as the effects of automated control vehicles [40], traffic smoothing strategies [81; 102], electric vehicles [53], and traffic incidents [103]. While generating more data is a crucial first step, the development of more analytical tools is equally paramount. The majority of existing tools do not open-source their analysis software (shown in Table 2), which potentially limits subsequent analyses on different datasets. This paper aims to fill this gap by exploring the analysis of massive trajectory data to make stop-and-go wave analysis at scale possible.
@@ 4 Methodology
$\bm{\mathbfit{\tau}}_{i}$ vector representing the trajectory of vehicle $i$ - -
$\bm{\mathbfit{\tau}}_{i}^{c}$ vector representing the stationary component trajectory of vehicle $i$ - -
$\bm{\mathbfit{\tau}}_{i}^{osc}$ vector representing the oscillation component trajectory of vehicle $i$ - -
@@ 4.1 Preliminaries
Long-range raw vehicle trajectory data as from [11] is imperfect due to the challenges of tracking and re-identification in dense traffic conditions [104; 105]. In essence, it is difficult to track vehicles perfectly across hundreds of cameras and through occlusions [106], so trajectory datasets over large spatial ranges often contain fragmented trajectories. These issues do not significantly impact local traffic measurements, such as traffic speeds, which can still be obtained with extremely high fidelity. We prepare raw data for the methods described next by utilizing the method from [107], which we describe in more detail in A. The resulting virtual trajectories have no passing or lane change maneuvers. This allows efficient search between adjacent trajectories, but the methods described next could also be applied to raw trajectory data with lane-change maneuvers with slight modifications.
@@ 4.2 Identifying the stop-and-go waves: the trajectory perspective
Let $\mathcal{T}$ be the set of all trajectories. Note that the preprocessing in A ensures that trajectories are ordered by increasing time, and no trajectories enter or exit the considered lane. Let $\bm{\mathbfit{\tau}}_{i}\in\mathcal{T}$ represent the trajectory of vehicle $i$ as a vector of time-space points. Throughout this section, we will use $\tau_{i}(t)$ to represent a continuous, functional representation of the trajectory of vehicle $i$ from which the points in $\bm{\mathbfit{\tau}}_{i}$ are sampled. To detect the critical events from the trajectories, each trajectory is decomposed into two components as follows:
@@
In our implementation, the local maxima and minima of $\tau_{i}^{\text{osc}}(t)$ are identified using peak-finding algorithms scipy.signal.find_peaks [108]. We find all local maxima and minima in $\bm{\mathbfit{\tau}}_{i}^{osc}$, yielding a set $d_{i}$ of wave front and a set $a_{i}$ of wave tail points for trajectory $i$. We keep only those wave fronts that are followed by a wave tail and those wave tails that are preceded by a wave front, excluding any wave fronts or tails that occur at the boundaries of the trajectories, as they do not contribute to a complete stop-and-go cycle. This ensures that there are the same number $K_{i}$ of wave front points and wave tail points for trajectory $i$. This process is repeated for all trajectories in $\mathcal{T}$, yielding an overall set of wave front and tail points $\mathcal{D}=\{d_{0},d_{1},...d_{i}\}$ and $\mathcal{A}=\{a_{0},a_{1},...a_{i}\}$.
@@ 5 Data
I-24 MOTION is a traffic instrument for the freeway traffic observation [112; 11; 105], which is designed for the continuous freeway traffic data collection and analysis. It is expected to generate 200 million vehicle miles traveled (VMT) in trajectory data annually [105], which poses significant challenges for data analysis. In this paper, the I-24 MOTION INCEPTION data released by [11] is used for the stop-and-go wave analysis with the proposed method in Section 4. Table 4 summarizes the data used in this paper, including the corresponding week day of the date, vehicle miles traveled (VMT) and the mean speed (MS). The method input consists of virtual trajectories generated from the mean speed field, designed to address data quality issues [104] in individual trajectories and enable scalable measurements since traffic waves are the phenomena in between microscopic and macroscopic. Further details are provided in A.
@@ 7.5 Spatial distribution of wave bifurcation and merge
Although identifying the root cause of bifurcation and merging is challenging, our analysis of road geometry suggests that these phenomena can occur anywhere, although with a concentration within the ramp impact area. We believe the underlying cause is more closely tied to driving behavior, with road geometry acting as a contributing factor by influencing the frequency of cut-ins and lane changes. Factors such as asymmetric driving behavior [115; 16], anticipation effects [116], and driver distraction [73] may play a significant role. A more detailed microscopic analysis of driving behavior, enabled by improvements in data quality [104], could provide deeper insights in the future.
@@ 8 Conclusion
The code developed in this article is publicly available at https://github.com/I24-MOTION/wave-analysis and will be listed on https://i24motion.org. The tools introduced in this article could open new possibilities for analyzing stop-and-go traffic waves at scale and it could lower the barrier for researchers to deal with large-scale trajectory data. For example, the I-24 MOTION instrument [11] will enable both CAV field experiments [81; 121] and ITS deployments [122]. It is worthy of attention that the method and tool presented in this article allow for empirically analyzing how control strategies and vehicle automation influence traffic wave properties, a growing topic within traffic flow theory and characteristics [123; 124; 125; 93; 120; 121]. Further research will focus on investigating the factors influencing wave evolution [65], accompanied by refinements in data quality.
@@ A.1 Virtual sensors and virtual vehicles
Long-range raw vehicle trajectory data is imperfect due to the challenges of tracking and re-identification in dense traffic conditions [104; 105]. The tasks fail when errors occurring at the current frame cannot be corrected using information from subsequent frames. These errors may arise from various system sources, including homography estimation [104] and object occlusion [106]. However, these challenges do not significantly impact local traffic measurements, such as traffic speeds, which can still be obtained with extremely high fidelity. In this context, generating vehicle trajectories [126] from high-fidelity speed field can serve as a surrogate measure for the raw trajectory data, enabling the analysis of stop-and-go waves.We use the virtual trajectory generation method from [107] as a first step in wave identification in this work. We detail this approach below. Table 3 provides a comprehensive list of the variables and parameters used throughout the appendix.
@@ A.1.1 Virtual sensors: Edie’s definition for macroscopic measurements
Edie [127] provides an approach to calculate spatio-temporal mean of density, flow speed, and traffic flow from vehicle trajectories, which is named as Edie’s definition (shown in Figure 15). According to the definition, the density and the flow can be computed from the total travel time (TTT) and total travel distance (TTD) of all the vehicles within an area. Using Edie’s definition allows for the parallel calculation of measurements on a trajectory-by-trajectory basis [107]. Consider a shear box [128; 126] centered at a point ($t$, $x$), and let $\Delta t$ and $\Delta x$ denote the height and width of the box, as shown in Figure 15. The macroscopic estimates can be computed as:
@@ A.1.2 Adaptive Smoothing
With the original speed field generated, virtual trajectories can be created. However, the original macroscopic measurements often contain missing values and outliers. As suggested by [126], an adaptive smoothing method should be applied to the raw data to improve accuracy. The Adaptive Smoothing Method (ASM), developed by [129; 130], is a widely used smoothing and interpolation algorithm for constructing a continuous spatio-temporal mean speed field. While it is especially useful for data collected from fixed infrastructure sensors, such as inductive loops, it is also applicable to the macroscopic speed data generated from trajectories. This method is particularly helpful in interpolating speeds in occluded areas, such as under bridges. We directly apply this method, with parameter settings as listed in Table 9.
@@
$\sigma$(mile) smoothing width in space coordinate 0.12
$\tau$(second) smoothing width in time coordinate 20